# fractional integral on Hardy spaces on product domains

### v2

@@ Abstract.
By using the vector-valued theory of singular integrals, we prove a Hardy–Littlewood–Sobolev inequality on product Hardy spaces $H^{p}_{\rm{prod}}$, which is a parallel result of the classical Hardy–Littlewood–Sobolev inequality. The same technique shows the $H^{p}_{\rm{prod}}$-boundedness of the iterated Hilbert transform. As a byproduct, new proofs of several recently discovered Hardy type inequalities on product Hardy spaces are obtained, which avoid complicated Calderón–Zygmund theory on product domain, rendering them considerably simpler than the original proofs.
@@ 1.1. The Hardy–Littlewood–Sobolev inequality
as long as $0<p<q<\infty$. This is proved by Krantz in 1982 ([13], Corollary 2.3) by using atomic decomposition of $H^{p}$.
@@ 1.2. Hardy space on product domain
We want to know if there exists a Hardy–Littlewood–Sobolev inequalty on $H^{p}_{\mathrm{prod}}(\mathbb{R}^{d})$. In the definition (1.1), there is only one parameter $\delta>0$, so the original $I_{\alpha}$ requires certain modification to coherent with the multi-parameter nature of $H^{p}_{\mathrm{prod}}(\mathbb{R}^{d})$. A nature candidate is the following operator:
@@
In Krantz’s proof of (1.2), the following key property of the ball is used: Let $B$ be a Euclidean ball centered at 0, if $x\in B$ and $y\notin 2B$, then $|x-y|\approx|y|$. The geometry of a general open set is far more complicated than balls, so no such property is available, and a step-by-step repetition of Krantz’s proof does not work.
@@
Nevertheless, there are still ways to circumvent these difficulties. In this paper, we show that, if we examine $H^{p}_{\rm{prod}}$ from the vector-valued point of view, then the $H^{p}_{\rm{prod}}\mapsto H^{q}_{\rm{prod}}$ boundedness can be easily established. Moreover, we do not need to resort any deep theory of $H^{p}_{\rm{prod}}$ (except its definition).
@@
Also, based on the same technique, new easier (both conceptually and technically) proofs of the main results of a recent paper [5] by Dyachenko, Nursultanov, Tikhonov and Weisz are obtained. An example is the $H^{p}_{\rm{prod}}$ Hardy–Littlewood inequality
@@ 3. Fractional integral on product spaces
In this section, we prove Theorem 1.1. We only consider the case $d=2$ for simplicity. Although theories of $H^{p}_{\rm{prod}}(\mathbb{R}^{2})$ and $H^{p}_{\rm{prod}}(\mathbb{R}^{3})$ are essentially different, for our approach used in this paper, the dimension $d$ does not cause problems. Also, the 1-dimensional case is known, as there is no product theory when $d=1$. The method of the proof goes back at least as far as [7] (see the lemma on Page 139 there). Write
@@
Recall that our goal is to calculate $\iint|S(I_{(\alpha,2)}f)|^{q}\,\mathrm{d}y\mathrm{d}x$, let us first calculate the $\int\mathrm{d}y$ integral:
@@ 4. New proof of several Hardy’s inequalities on product spaces
This criterion is due to Fefferman (see [6], page 840). A direct generalization to $H^{1}_{\rm{prod}}(\mathbb{R}^{3})$ is not true, as there exists a counter-example due to Journé (see Section 2 of [12]). To apply this criterion to $H^{p}_{\rm{prod}}(\mathbb{R}^{d})$ for $d\geq 3$ and $0<p\leq 1$, the authors in [5] used a variant form of it, see Lemma 6 there (without proofs), or [20], Chapter 3, Theorem 3.6.12. The proof of Theorem 3.6.12 in [20] (takes nearly 10 pages), along with some newly introduced definitions (like the simple $p$-atoms), requires considerable work.
Worse then, Theorem 3.6.12 and Fefferman’s criterion are results for Lebesgue measure. Due to the nature of the problem here, in Lemma 6 of [5], the authors apply a (further) variant: replacing the Lebesgue measure in the left hand side of (4.2) by $\mathrm{d}\eta\coloneqq|x_{1}x_{2}\cdots x_{d}|^{-2}\mathrm{d}x$. Therefore, an additional verification of a majorant property of $\mathrm{d}\eta$ is required. The reader may find this property in [5], inequality (38).
The rest of the job is straightforward: Check the decay condition (4.2) by subtracting certain Taylor polynomial of higher degree, and then conclude the proof. However, the calculation of this verification is quite lengthy.
@@
The authors in [5] choose another operator $T_{\mathcal{H}}(f)(\xi)\coloneqq\Big(\prod_{i=1}^{d}\xi_{i}\Big)\mathcal{H}(\widehat{f})(\xi)$, and then proceed as in the proof of Theorem 4.1. Although these proofs are similar, since $T$ (see (4.1) for its definition) and $T_{\mathcal{H}}$ have different form, there seems no way to be lazy, and one has repeat the long steps in the proof of Theorem 4.1 faithfully.
@@
As a byproduct of Proposition 3.1, we give new easier proofs of several main results in [5], which does not rely on complicated Calderón–Zygmund theory on product domain. The following claims will be used.
@@
By definition of $H^{p}$ and $f\in L^{2}$, the Poisson maximal function $\sup_{\delta>0}|f\ast P_{\delta}|\in L^{2}$. We then subtract an a.e. convergent sequence and use Fatou’s lemma. Same for $H^{p}_{\rm{prod}}$ spaces, see [9], proofs of Theorem  1.1 and Theorem 1.2.
@@
$\|f\|_{H^{p}(\mathbb{R})}\approx\|f\|_{L^{p}(\mathbb{R})}+\|Hf\|_{L^{p}(\mathbb{R})}.$
In fact, [19] states that $\|f\|_{H^{p}(\mathbb{R})}\lesssim\|f\|_{L^{p}(\mathbb{R})}+\|Hf\|_{L^{p}(\mathbb{R})}$, but by Proposition 4.4, we know that $\|f\|_{L^{p}(\mathbb{R})}\lesssim\|f\|_{H^{p}(\mathbb{R})},$ and $\|Hf\|_{L^{p}(\mathbb{R})}\lesssim\|f\|_{H^{p}(\mathbb{R})}$. We need $f\in L^{2}(\mathbb{R})$ to ensure that (by Proposition 4.4) the quantity $\|f\|_{L^{p}(\mathbb{R})}$ is well-defined.
@@
$\int_{\mathbb{R}^{2}}\frac{|\widehat{f}(t_{1},t_{2})|^{p}}{|t_{1}|^{2-p}|t_{2}|^{2-p}}\,\mathrm{d}t\lesssim\|f\|^{p}_{p}+\|H_{1}f\|^{p}_{p}+\|H_{2}f\|^{p}_{p}+\|H_{1}H_{2}f\|^{p}_{p},.$
@@
In order to use Proposition 4.5, we employed the following result: For a.e. $t_{1},t_{2}\in\mathbb{R}$, the function
(4.3) $z\mapsto\mathcal{F}_{2}\mathcal{F}_{1}f(t_{1},t_{2},z)=\int_{\mathbb{R}^{2}}f(x,y,z)\mathrm{e}^{-2\pi\mathrm{i}(xt_{1}+yt_{2})}\,\mathrm{d}x\mathrm{d}y\in L^{2}(\mathbb{R},\mathrm{d}z).$
The equality “==” in (4.3) holds beacuse $f\in L^{1}$. To prove the $L^{2}(\mathbb{R},\mathrm{d}z)$ integrability, just notice that for a.e. $t_{1},t_{2}$,
$\|\mathcal{F}_{2}\mathcal{F}_{1}f(t_{1},t_{2},z)\|^{2}_{L^{2}(\mathbb{R},\mathrm{d}z)}=\|\mathcal{F}_{3}\mathcal{F}_{2}\mathcal{F}_{1}f(t_{1},t_{2},t_{3})\|^{2}_{L^{2}(\mathbb{R},\mathrm{d}t_{3})}\leq\|f(x,y,z)\|^{2}_{L^{2}(\mathbb{R}^{3})}.$
@@
$\int_{\mathbb{R}}\frac{1}{|x_{1}|^{2}}\int_{\mathbb{R}}\frac{1}{|x_{2}|^{2}}\bigg|\int_{0}^{x_{1}}\int_{0}^{x_{2}}\widehat{f}(t)\,\mathrm{d}t_{1}\mathrm{d}t_{2}\bigg|^{p}\,\mathrm{d}x_{1}\mathrm{d}x_{2}$
By definition of Fourier transform and Fubini’s theorem, we write the innermost integral $\int_{0}^{x_{1}}\int_{0}^{x_{2}}$ as
@@
where $f_{x_{1}}(y_{2})\coloneqq\int_{0}^{x_{1}}\int_{\mathbb{R}}f(y_{1},y_{2})\mathrm{e}^{-2\pi\mathrm{i}y_{1}t_{1}}\,\mathrm{d}y_{1}\mathrm{d}t_{1}$. By the one-dimensional result [^5] According to [5], for this theorem, even the 1-dimensional case is a new result. However, the 1-dimensional case can be easily proved by atomic decomposition, so we omit it here and admit it as true.
@@
It is easy to show that for almost everywhere $x_{1}\in\mathbb{R}$, the function $y_{2}\mapsto f_{x_{1}}(y_{2})$ belongs to $L^{2}(\mathbb{R},\mathrm{d}y_{2})$. So $\|f_{x_{1}}\|^{p}_{H^{p}(\mathbb{R},\mathrm{d}y_{2})}\approx\|f_{x_{1}}\|^{p}_{L^{p}(\mathbb{R},\mathrm{d}y_{2})}+\|H_{2}f_{x_{1}}\|^{p}_{L^{p}(\mathbb{R},\mathrm{d}y_{2})}$. By previous estimate, we have
@@
Again, the 1-dimensional result gives
@@
where $f_{x_{1}}(t_{2})\coloneqq\int_{0}^{x_{1}}f(t_{1},t_{2})\,\mathrm{d}t_{1}$. The 1-dimensional inequality [^6] According to [5], the 1-dimensional inequality is classical, so we use it freely here. states that

### v3

@@ Abstract.
Using the vector-valued theory of singular integrals, we prove a Hardy–Littlewood–Sobolev inequality on product Hardy spaces $H^{p}_{\rm{prod}}$, which is a parallel result of the classical Hardy–Littlewood–Sobolev inequality. The same technique shows the $H^{p}_{\rm{prod}}$-boundedness of the iterated Hilbert transform. As a byproduct, shorter proofs of several recently discovered Hardy type inequalities on product Hardy spaces are obtained.
@@ 1.1. The Hardy–Littlewood–Sobolev inequality
as long as $0<p<q<\infty$. This is proved by Krantz in 1982 ([13], Corollary 2.3) using atomic decomposition of $H^{p}$.
@@ 1.2. Hardy space on product domain
We want to know if there exists a Hardy–Littlewood–Sobolev inequalty on $H^{p}_{\mathrm{prod}}(\mathbb{R}^{d})$. In the definition (1.1), there is only one parameter $\delta>0$, so the original operator $I_{\alpha}$ requires certain modification to coherent with the multi-parameter nature of $H^{p}_{\mathrm{prod}}(\mathbb{R}^{d})$. A nature candidate is the following operator:
@@
In Krantz’s proof of (1.2), the following geometric property is used: Let $B$ be a Euclidean ball centered at 0, if $x\in B$ and $y\notin 2B$, then $|x-y|\approx|y|$. The geometry of a general open set is far more complicated than balls, so no such property is available, and a step-by-step repetition of Krantz’s proof does not work.
@@
Nevertheless, there are still ways to circumvent these difficulties. In this paper, we show that, if we examine $H^{p}_{\rm{prod}}$ from the vector-valued point of view, then the $H^{p}_{\rm{prod}}\mapsto H^{q}_{\rm{prod}}$ boundedness can be easily established. Moreover, we do not need to resort deep results of $H^{p}_{\rm{prod}}$.
@@
Also, based on the same technique, we obtain shorter (both conceptually and technically) proofs of the main results of a recent paper [5] by Dyachenko, Nursultanov, Tikhonov and Weisz. An example is the $H^{p}_{\rm{prod}}$ Hardy–Littlewood inequality
@@ 3. Fractional integral on product spaces
In this section, we prove Theorem 1.1. We only consider the case $d=2$ for simplicity. Although theories of $H^{p}_{\rm{prod}}(\mathbb{R}^{2})$ and $H^{p}_{\rm{prod}}(\mathbb{R}^{3})$ are essentially different, for our approach used in this paper, the dimension does not cause problems. Also, the 1-dimensional case is known, as there is no product theory when $d=1$. The method of the proof goes back at least as far as [6] (see the lemma on Page 139 there). Write
@@
Our goal is to calculate $\iint|S(I_{(\alpha,2)}f)|^{q}\,\mathrm{d}y\mathrm{d}x$, let us first calculate the $\int\mathrm{d}y$ integral:
@@ 4. New proof of several Hardy’s inequalities on product spaces
This criterion is due to Fefferman (see [7], page 840). A direct generalization to $H^{1}_{\rm{prod}}(\mathbb{R}^{3})$ is not true, as there exists a counter-example due to Journé (Section 2 in [12]). To apply this criterion to $H^{p}_{\rm{prod}}(\mathbb{R}^{d})$ for $d\geq 3$ and $0<p\leq 1$, the authors in [5] used a variant form, see Lemma 6 there, or [21], Chapter 3, Theorem 3.6.12. The proof of Theorem 3.6.12 in [21], along with some newly introduced definitions (like the simple $p$-atoms), requires some preparatory work.
Moreover, Fefferman’s criterion is a result for Lebesgue measure. Due to the nature of the problem here, in Lemma 6 of [5], the authors apply a (further) variant: replacing the Lebesgue measure in the left hand side of (4.2) by $\mathrm{d}\eta\coloneqq|x_{1}x_{2}\cdots x_{d}|^{-2}\mathrm{d}x$. Therefore, an additional verification of a majorant property (see inequality (38) in [5]) of $\mathrm{d}\eta$ is required.
@@
To prove this theorem, the authors in [5] define an operator $T_{\mathcal{H}}(f)(\xi)\coloneqq\Big(\prod_{i=1}^{d}\xi_{i}\Big)\mathcal{H}(\widehat{f})(\xi)$, and then proceed as in the proof of Theorem 4.1.
@@
As a byproduct of Proposition 3.1, we give shorter proofs of several main results in [5], which does not rely on Calderón–Zygmund theory on product domain. The following claims will be used.
@@
Since $f\in L^{2}$, the Poisson maximal function $\sup_{\delta>0}|f\ast P_{\delta}|\in L^{2}$. We then subtract an a.e. convergent sequence and use Fatou’s lemma to obtain $L^{p}$-integrability of $f$. For details, see [9], proofs of Theorem  1.1 and Theorem 1.2.
@@
$\|f\|_{H^{p}(\mathbb{R})}\approx\|f\|_{L^{p}(\mathbb{R})}+\|Hf\|_{L^{p}(\mathbb{R})},$
where $H$ is the Hilbert transform.
The monogrpah [19] states that $\|f\|_{H^{p}(\mathbb{R})}\lesssim\|f\|_{L^{p}(\mathbb{R})}+\|Hf\|_{L^{p}(\mathbb{R})}$, but by Proposition 4.4, we know that $\|f\|_{L^{p}(\mathbb{R})}\lesssim\|f\|_{H^{p}(\mathbb{R})},$ and $\|Hf\|_{L^{p}(\mathbb{R})}\lesssim\|f\|_{H^{p}(\mathbb{R})}$. The $L^{2}(\mathbb{R})$-integrability of $f$ ensures that (by Proposition 4.4) the quantity $\|f\|_{L^{p}(\mathbb{R})}$ is well-defined.
@@
$\int_{\mathbb{R}^{2}}\frac{|\widehat{f}(t_{1},t_{2})|^{p}}{|t_{1}|^{2-p}|t_{2}|^{2-p}}\,\mathrm{d}t\lesssim\|f\|^{p}_{p}+\|H_{1}f\|^{p}_{p}+\|H_{2}f\|^{p}_{p}+\|H_{1}H_{2}f\|^{p}_{p}.$
@@
In order to use Proposition 4.5, we employed the following corollary of Fubini’s theorem and Plancherel’s identity: For a.e. $t_{1},t_{2}\in\mathbb{R}$, the following function
(4.3) $z\mapsto\mathcal{F}_{2}\mathcal{F}_{1}f(t_{1},t_{2},z)=\int_{\mathbb{R}^{2}}f(x,y,z)\mathrm{e}^{-2\pi\mathrm{i}(xt_{1}+yt_{2})}\,\mathrm{d}x\mathrm{d}y$
is $L^{2}(\mathbb{R},\mathrm{d}z)$-integrable.
@@
$\int_{\mathbb{R}}\frac{1}{|x_{1}|^{2}}\int_{\mathbb{R}}\frac{1}{|x_{2}|^{2}}\bigg|\int_{0}^{x_{1}}\int_{0}^{x_{2}}\widehat{f}(t)\,\mathrm{d}t_{1}\mathrm{d}t_{2}\bigg|^{p}\,\mathrm{d}x_{1}\mathrm{d}x_{2}.$
We write the innermost integral $\int_{0}^{x_{1}}\int_{0}^{x_{2}}$ as
@@
where $f_{x_{1}}(y_{2})\coloneqq\int_{0}^{x_{1}}\int_{\mathbb{R}}f(y_{1},y_{2})\mathrm{e}^{-2\pi\mathrm{i}y_{1}t_{1}}\,\mathrm{d}y_{1}\mathrm{d}t_{1}$. By one-dimensional result [^5] According to [5], for this theorem, even the 1-dimensional case is a new result. However, the 1-dimensional case can be easily proved by atomic decomposition, so we omit it here and admit it as true.
@@
It is easy to show that for almost everywhere $x_{1}\in\mathbb{R}$, the function $y_{2}\mapsto f_{x_{1}}(y_{2})$ belongs to $L^{2}(\mathbb{R},\mathrm{d}y_{2})$, so $\|f_{x_{1}}\|^{p}_{H^{p}(\mathbb{R},\mathrm{d}y_{2})}\approx\|f_{x_{1}}\|^{p}_{L^{p}(\mathbb{R},\mathrm{d}y_{2})}+\|H_{2}f_{x_{1}}\|^{p}_{L^{p}(\mathbb{R},\mathrm{d}y_{2})}$. By previous estimate, we have
@@
Again, one-dimensional result gives
@@
where $f_{x_{1}}(t_{2})\coloneqq\int_{0}^{x_{1}}f(t_{1},t_{2})\,\mathrm{d}t_{1}$. The one-dimensional inequality [^6] According to [5], the 1-dimensional inequality is classical, so we use it freely here. states that
@@
5. Acknowledgement
This paper was completed at Nicolaus Copernicus University in Toruń, where the author was a postdoctoral researcher at that time. The author would like to thank Yuri Tomilov for introducing him to the theory of Hardy spaces and Hardy type inequalities.